it is known that the sg (specific gravity) of oil is 0.95 sg of mercury (hg) is 13.6. h1 = 90cm h2 = 15 cm h3 = 50 cm
1
Expert's answer
2020-12-07T03:46:33-0500

We know that ,

specific gravity is related with density as

density of substance= specific gravity "\times" density of water

here ,

and pressure relation with density is

"P_1=\rho gh_1" so from this relation we can find pressure and density of substance

density of substance ="0.95\times 1000=950 kg/m^3,"


"P_1=950\times 9.8\times" 0.9=8379 Pa